OpenStudy (anonymous):

Jeannie is 8 years older than her younger sister April. When added together, their ages equal 198. How old is April?

OpenStudy (zzr0ck3r):

jeannie - 8 years =age of april so j - 8 = a jeannies age + Aprils age = 198 so j+a=198 so we have two equations 1) j+a=198 and 2) j-8=a so plug the 2nd into the first for a j+a=j+j-8=198 2j=206 j=103 so a=j-8 = 103-8 = 95 april is 95

OpenStudy (anonymous):

any time you have a problem like this with ages, subtract the difference and divide by 2 to get the lower of the 2 ages 198 - 8 = 190 190/2 = 95

OpenStudy (anonymous):

95